# arXiv:2603.03251v1[cs.LG] 03 Mar 2026

### v1

@@ Abstract
Autoregressive decoding is bottlenecked by its sequential nature. Speculative decoding has become a standard way to accelerate inference by using a fast draft model to predict upcoming tokens from a slower target model, and then verifying them in parallel with a single target model forward pass. However, speculative decoding itself relies on a sequential dependence between speculation and verification. We introduce speculative speculative decoding (SSD) to parallelize these operations. While a verification is ongoing, the draft model predicts likely verification outcomes and prepares speculations pre-emptively for them. If the actual verification outcome is then in the predicted set, a speculation can be returned immediately, eliminating drafting overhead entirely. We identify three key challenges presented by speculative speculative decoding, and suggest principled methods to solve each. The result is Saguaro, an optimized SSD algorithm. Our implementation is up to 2x faster than optimized speculative decoding baselines and up to 5x faster than autoregressive decoding with open source inference engines.
@@ 1 Introduction
Altogether, Saguaro achieves up to 2x speedup over optimized speculative decoding and up to 5x over autoregressive generation (Figure 7), while strictly improving the throughput-latency Pareto frontier across a range of batch sizes.
@@ 2.2 Related Work
Parallel Speculative Decoding. AMUSD (McDanel, 2025) and PEARL (Liu et al., 2025) propose speculating the next round during ongoing verification, but only prepare for the verification outcome in which all tokens are accepted, which is one special case of many possible outcomes. SwiftSpec (Zhang et al., 2025) and SpecBranch (Shen et al., 2025) prepare a larger cache consisting of a token tree branching off of the speculation being verified (thus enabling larger speedups), but both use fallback strategies that do not work at large batch sizes. SpecBranch (Shen et al., 2025) has motivation similar to SSD’s: it constitutes (approximately) a special case of the SSD framework where only a single branching point is allowed, where the fallback speculator is equal to the regular speculator, and where the hyperparameters (branching point, number branches, speculation length) are dynamically chosen to maximize speedups. SwiftSpec (Zhang et al., 2025) considers the special case of greedy sampling, and adopts a fallback strategy of just-in-time speculation that struggles at higher temperatures and batch sizes when cache-misses become inevitable.
@@
Tree-Based Speculative Decoding. Numerous speculative decoding methods have been proposed that increase the expected number of accepted tokens by allowing the draft model to speculate a tree of tokens instead of a sequence, thereby giving the verifier several token options at each position (Miao et al., 2024; Li et al., 2024b; Chen et al., 2024; Svirschevski et al., 2024). Our method differs in important ways: First, existing tree-based methods are still sequential: speculate, then verify. Second, tree-based methods introduce a large amount of verifier compute—which is quite expensive due to the size of the target model—because the entire tree must be verified in the target model forward pass. Our method, on the other hand, scales up the speculation compute by pre-speculating for many verification outcomes in parallel, but does not introduce any additional verifier compute. Lastly, it is important to note that our method can be combined with these tree-based approaches for further gains (see Appendix E for discussion).
@@
Improved Draft Architectures. There have been many advancements in draft model architectures that can improve the acceptance rates and/or speeds of the draft model and which can be used in SSD for even better performance. For example, EAGLE (Li et al., 2024a, b, 2025b) allows the draft model to take as input the powerful representations from the target model for the current prefix, while GliDe (Du et al., 2024) and LongSpec (Yang et al., 2025) allow the draft model to perform cross-attention on the KV cache of the target model. Alternate model architectures, like diffusion LLMs (Nie et al., 2025; Sahoo et al., 2024; Li et al., 2025a; Christopher et al., 2025; Samragh et al., 2025; Chen et al., 2026) or SSMs (Gu et al., 2022; Gu & Dao, 2023; Wang et al., 2024), can also be used to increase the speed with which the draft model can produce the speculated token sequence. We include more technical details on how SSD can fruitfully be combined with these improved draft model architectures (e.g., EAGLE-3; Li et al. (2025b)) in Appendix E.
@@ 3.1 Algorithm
For intuition on the SSD algorithm and why it is lossless, consider the similarities with speculative execution (González & González (1997)) for CPU optimization. Like speculative execution—which uses idle compute to pre-execute conditional paths in case they are later needed—SSD pre-speculates over many possible verification outcomes so that, when one occurs, the corresponding token sequence is immediately available. That sequence is then verified exactly as in SD. If the realized outcome was not pre-speculated, SSD falls back to a synchronous speculator, which is lossless since it reduces to ordinary SD. The following definitions will be important in understanding the SSD framework.
@@ 4.3.3 Empirical Evaluation
We show in Figure 6 (left) that as the batch size increases, using a fast backup speculator that returns random tokens outperforms using a slow but more accurate neural speculator just-in-time. We note that for all of these batch sizes, we use fan-out $F=1$ and 1 H100 GPU for drafting. In Figure 6 (right) we show that we can reduce latency further by scaling draft-time compute. Specifically, we show the projected speedups we would get at batch size 16—computed based on our cache hit rate curves (Figure 4)—if we increase the number of draft GPUs (and thus, the fan-out factor). Dedicating more devices to speculation leads to larger speculation caches and thus higher cache hit rates.
@@ 5 End-to-End Evaluation
In Figure 7 we compare the end-to-end performance of Saguaro to key baselines, including our own implementation of ordinary speculative decoding, as well as the strongest open source inference engines (vLLM/SGLang). [^3] We find SGLang usually outperforms vLLM, so we use it as the main baseline throughout. We also include EAGLE-3 when supported by open source engines. [^4] In our experiments Llama-3.2-1B outperforms Eagle-3 as a draft model for Llama-3.1-70B in SGLang. We attain almost a 5x speedup on average compared to our autoregressive baseline, and are up to twice as fast as the optimized speculative decoding baselines, establishing a new state of the art.
@@ 6 Conclusion and Limitations
Speculative decoding trades compute for latency, but drafting and verification remain synchronous. We introduce speculative speculative decoding (SSD), which parallelizes even this sequential dependence. We derive performance bounds, study each component of the design space, and distill our findings into Saguaro, an optimized SSD algorithm which achieves up to 2x speedup over optimized speculative decoding and up to 5x over autoregressive generation across datasets and model families.
@@ Acknowledgements
We gratefully acknowledge the support of the Schmidt Sciences AI2050 fellowship, the Google ML and Systems Junior Faculty Awards, and the Google Research Scholar program. We thank Together AI for supporting this project with compute resources. Tanishq did this work during an internship at Together AI, and thanks Tatsunori Hashimoto and Luke Bailey for helpful comments and discussion.
@@ Appendix E Combining SSD with SD Variants
Token-Tree Methods. Token-tree SD methods (Miao et al., 2024; Chen et al., 2024; Li et al., 2024b, 2025b) can be combined naturally with SSD. SSD drafts a linear chain for many possible verification outcomes. To combine SSD with these token-tree approaches, one would simply pre-speculate (and verify) a token tree for each verification outcome instead of a token chain. In practice, this requires a more elaborate attention mask than that presented in Figure 8.

### v3

# Speculative Speculative Decoding

@@ Abstract
Autoregressive decoding is bottlenecked by its sequential nature. Speculative decoding has become a standard way to accelerate inference by using a fast draft model to predict upcoming tokens from a slower target model, and then verifying them in parallel with a single target model forward pass. However, speculative decoding itself relies on a sequential dependence between speculation and verification. We introduce speculative speculative decoding (SSD) to parallelize these operations. While a verification is ongoing, the draft model predicts likely verification outcomes and prepares speculations pre-emptively for them. If the actual verification outcome is in the predicted set, a speculation can be returned immediately, eliminating drafting overhead entirely. We identify three key challenges presented by speculative speculative decoding, and suggest principled methods to solve each. The result is Saguaro, an optimized SSD algorithm. Our implementation is on average 30% faster than the strongest speculative decoding baselines and up to 5x faster than autoregressive decoding with open source inference engines.
@@ 1 Introduction
Altogether, Saguaro is on average 30% faster than the strongest speculative decoding baselines and up to 5x faster than autoregressive generation (Figure 7), while strictly improving the throughput-latency Pareto frontier across a range of batch sizes.
@@ 2.2 Related Work
Parallel Speculative Decoding. AMUSD (McDanel, 2025) and PEARL (Liu et al., 2025) propose speculating the next round during ongoing verification, but only prepare for the verification outcome in which all tokens are accepted, which is one special case of many possible outcomes. DSI (Timor et al., 2025) also overlaps drafting and verification, verifying multiple sequences in parallel with distinct logical copies of the verifier model. Mirror-SD (Bhendawade et al., 2025) includings branch-complete speculative rollouts conditioned on early-exit signals from the target model and heterogeneous execution across devices. SwiftSpec (Zhang et al., 2025) and SpecBranch (Shen et al., 2025) prepare a larger cache consisting of a token tree branching off of the speculation being verified (thus enabling larger speedups), but both use fallback strategies that do not work at large batch sizes. SpecBranch (Shen et al., 2025) constitutes (approximately) a special case of the SSD framework where only a single branching point is allowed and where the fallback speculator is equal to the regular speculator. SwiftSpec (Zhang et al., 2025) considers the special case of greedy sampling, and adopts a fallback strategy of just-in-time speculation that struggles at higher temperatures and batch sizes when cache-misses become inevitable.
@@
Tree-Based Speculative Decoding. Numerous speculative decoding methods have been proposed that increase the expected number of accepted tokens by allowing the draft model to speculate a tree of tokens instead of a sequence, thereby giving the verifier several token options at each position (Miao et al., 2024; Li et al., 2024a; Chen et al., 2024; Svirschevski et al., 2024). Our method differs in important ways: First, existing tree-based methods are still sequential: speculate, then verify. Second, tree-based methods introduce a large amount of verifier compute—which is quite expensive due to the size of the target model—because the entire tree must be verified in the target model forward pass. Our method, on the other hand, scales up the speculation compute by pre-speculating for many verification outcomes in parallel, but does not introduce any additional verifier compute. Lastly, it is important to note that our method can be combined with these tree-based approaches for further gains (see Appendix E for discussion).
@@
Improved Draft Architectures. There have been many advancements in draft model architectures that can improve the acceptance rates and/or speeds of the draft model and which can be used in SSD for even better performance. For example, EAGLE (Li et al., 2024b, a, 2025b) allows the draft model to take as input the powerful representations from the target model for the current prefix, while GliDe (Du et al., 2024) and LongSpec (Yang et al., 2025) allow the draft model to perform cross-attention on the KV cache of the target model. Alternate model architectures, like diffusion LLMs (Nie et al., 2025; Sahoo et al., 2024; Li et al., 2025a; Christopher et al., 2025; Samragh et al., 2025; Chen et al., 2026) or SSMs (Gu et al., 2022; Gu and Dao, 2023; Wang et al., 2024), can also be used to increase the speed with which the draft model can produce the speculated token sequence. We include more technical details on how SSD can fruitfully be combined with these improved draft model architectures (e.g., EAGLE-3; Li et al. (2025b)) in Appendix E.
@@ 3.1 Algorithm
For intuition on the SSD algorithm and why it is lossless, consider the similarities with speculative execution (González and González (1997)) for CPU optimization. Like speculative execution—which uses idle compute to pre-execute conditional paths in case they are later needed—SSD pre-speculates over many possible verification outcomes so that, when one occurs, the corresponding token sequence is immediately available. That sequence is then verified exactly as in SD. If the realized outcome was not pre-speculated, SSD falls back to a synchronous speculator, which is lossless since it reduces to ordinary SD. The following definitions will be important in understanding the SSD framework.
@@ 4.3.3 Empirical Evaluation
We show in Figure 6 (left) that as the batch size increases, using a fast backup speculator that returns random tokens outperforms using a slow but more accurate neural speculator just-in-time. In Figure 6 (right) we show that we can reduce latency further by scaling draft-time compute. Specifically, we show the projected speedups we would get at batch size 16—computed based on our cache hit rate curves (Figure 4)—if we increase the number of draft GPUs (and thus, the fan-out factor). Dedicating more devices to speculation leads to larger speculation caches and thus higher cache hit rates.
@@ 5 End-to-End Evaluation
In Figure 7 we compare the end-to-end performance of Saguaro to key baselines, including our own implementation of ordinary speculative decoding, as well as the strongest open source inference engines (vLLM/SGLang). [^3] We find SGLang usually outperforms vLLM, so we use it as the main baseline throughout. We also include EAGLE-3 when supported by open source engines. [^4] In our experiments Llama-3.2-1B outperforms Eagle-3 as a draft model for Llama-3.1-70B in SGLang. We attain almost a 5x speedup on average compared to our autoregressive baseline, and are on average 30% faster than the strongest speculative decoding baselines, establishing a new state of the art.
@@ 6 Conclusion and Limitations
Speculative decoding trades compute for latency, but drafting and verification remain synchronous. We introduce speculative speculative decoding (SSD), which parallelizes even this sequential dependence. We derive performance bounds, study each component of the design space, and distill our findings into Saguaro, an optimized SSD algorithm which is on average 30% faster than the strongest speculative decoding baselines and up to 5x faster than autoregressive generation across datasets and model families.
@@ Acknowledgements
We gratefully acknowledge the support of the Schmidt Sciences AI2050 fellowship, the Google ML and Systems Junior Faculty Awards, and the Google Research Scholar program. We thank Together AI for supporting this project with compute resources. Tanishq thanks Tatsunori Hashimoto and Luke Bailey for helpful comments and discussion.
@@ B.3 Numerical Results
B.4 Baseline Sweeps
Cells report decode throughput (tok/s); $\dagger$ indicates incompatible configs due to known bugs in the library version used (SGLang 0.5.9 / vLLM 0.16.0). The best configuration per method is bolded. We sweep speculative lookahead and tree size for both vanilla SD and EAGLE-3 for SGLang, and just lookahead for vLLM since tree verification codepaths are known to be incomplete.
SGLang EAGLE-3 (steps $\times$ topk)
steps $\backslash$ topk 1 2 3 4 5
2 122.8 124.7 127.3 $\dagger$ $\dagger$
4 157.1 157.8 164.9 $\dagger$ $\dagger$
6 171.3 178.9 183.0 $\dagger$ $\dagger$
8 172.4 183.9 192.8 190.2 $\dagger$
10 168.7 183.0 192.5 $\dagger$ 186.9
12 164.8 178.7 186.3 187.6 185.4
14 158.4 173.8 180.3 182.9 176.9
SGLang Standalone (steps $\times$ topk)
steps $\backslash$ topk 1 2 3 4 5
2 123.0 123.9 125.4 $\dagger$ $\dagger$
4 159.4 161.2 164.9 $\dagger$ $\dagger$
6 176.7 182.3 186.7 $\dagger$ 164.9
8 182.1 192.3 197.7 $\dagger$ 163.1
10 181.5 195.7 201.9 178.0 160.3
12 180.4 195.2 199.0 $\dagger$ $\dagger$
14 175.7 193.6 $\dagger$ 171.5 $\dagger$
Table 1: SGLang 0.5.9 Llama-3.1-70B/1B baseline sweeps (decode tok/s).
Method $K=2$ $K=4$ $K=6$ $K=8$ $K=10$ $K=12$
vLLM EAGLE-3 143.7 181.4 197.6 197.0 191.8 187.9
vLLM Standalone 136.0 117.9 100.9 89.1 78.5 70.5
Table 2: vLLM 0.16.0 Llama-3.1-70B/1B baseline sweeps over speculative lookahead $K$ (decode tok/s).
@@ Appendix E Combining SSD with SD Variants
Token-Tree Methods. Token-tree SD methods (Miao et al., 2024; Chen et al., 2024; Li et al., 2024a, 2025b) can be combined naturally with SSD. SSD drafts a linear chain for many possible verification outcomes. To combine SSD with these token-tree approaches, one would simply pre-speculate (and verify) a token tree for each verification outcome instead of a token chain. In practice, this requires a more elaborate attention mask than that presented in Figure 8.